3. The time for a swing to move forward and backward can be determined by the formula t=2 pie square toot l/9.8 T represents the time, in seconds, taken by the swing to move through one complete cycle (forward and back) and L represents the length of the rope supporting the swing.

Determine the length of the rope supporting a swing that takes 3.4 seconds to move through one complete cycle. Show all steps and round the answer to the nearest tenth of a metre.

Answer by Boreal(15235)

Time=2*pi*sqrt(L/9.8)
3.4=2*pi(sqrt(L/9.8)
0.5411=sqrt (L/9.8), after dividing by 2pi
square both sides
0.2928=L/9.8
L=2.9 meters (not rounding intermediate calculations and rounding only at the end)